Write an equation in slope-intercept form of the line with the given properties.
Passes through (−2, 2) and (3, −3) y=

Answer by harpazo(655)

Find slope.
Let m = slope

m = -3-2/3-(-2)
m = -5/5 = -1
Use point-slope formula and one of the given points and our slope just found.

y - (-3) = -(x-3)
y + 3 = -x + 3
y = -x +3-3

y = -x + 0
Done.